Convert to percent notation.
What percent of 125 is 30?

Answer by stanbon(75887)

Convert to percent notation.
What percent of 125 is 30?
---
x/100 = 30/125
---
x = 100(30/125)
---
x = 3000/125
---
x = 24
==============
Ans: 24%
=================